Established 95% and 99% confidence interval about the mean for a psychological test that is given to thirty two persons and has a mean of 53 and a standard deviation of 7.5 assuming that the population standard deviation is not known.


1
Expert's answer
2021-10-20T14:25:47-0400

1. 95% confidence interval 

The critical value for "\alpha = 0.05" and "df = n-1 = 31" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.039513"

The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-t_c\times \dfrac{s}{\sqrt{n}}, \bar{X}+t_c\times \dfrac{s}{\sqrt{n}})"

"=(53- 2.039513\times \dfrac{7.5}{\sqrt{32}}, 53+2.039513\times \dfrac{7.5}{\sqrt{32}})"

"=(50.296, 55.704)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "50.296 < \mu < 55.704," which indicates that we are 95% confident that the true population mean "\mu"

is contained by the interval "(50.296, 55.704)."



2. 99% confidence interval 

The critical value for "\alpha = 0.01" and "df = n-1 = 31" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.744042"

The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-t_c\times \dfrac{s}{\sqrt{n}}, \bar{X}+t_c\times \dfrac{s}{\sqrt{n}})"

"=(53- 2.744042\times \dfrac{7.5}{\sqrt{32}}, 53+2.744042\times \dfrac{7.5}{\sqrt{32}})"

"=(49.362, 56.638)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "49.362 < \mu < 56.638," which indicates that we are 99% confident that the true population mean "\mu"

is contained by the interval "(49.362, 56.638)."